# Mechanisms are Transferable: Data-Efficient Low-Resource Adaptation via Circuit-Targeted Supervised Fine-Tuning

### v1

@@ D.2 Selection-ratio sweep for circuit construction (2%, 5%, 10%)
D.3 Shared vs. Held-out Stage-1 Pool for Circuit Discovery (Iteration 0 Stability)
@@
Rather than constructing the baseline mean $\mu$ and the circuit-discovery inputs from fully disjoint data sources, we intentionally draw (i) the Stage-1 fine-tuning examples, (ii) the balanced-mean pool used to estimate $\mu$, and (iii) the circuit-discovery set from the same underlying training pool. This design choice is primarily practical and stability-driven: label-balanced mean estimation is constrained by the class distribution in low-resource splits, and using a separate, disjoint pool can reduce the availability of balanced samples and increase variance in $\mu$.
Importantly, this shared-pool choice is not merely about allowing overlap between the mean pool and discovery set; it is about ensuring that $\mu$ and the head-relevance estimates are computed from the same in-distribution pool as the checkpoint’s Stage-1 adaptation. This yields more stable circuits (especially at early iterations) and avoids degenerate selections that can arise when the baseline and discovery inputs are drawn from different pools. Concretely, ‘shared-pool’ refers to using Pool A for Stage-1 fine-tuning, µ estimation, and CD-T discovery; Pool B remains disjoint and is used only for Stage-2 tuning.
We observe in our low-resource setting that CD-T’s iteration-0 selection is more stable when circuit discovery is performed on the same pool used for Stage-1 fine-tuning, rather than on a held-out subset from that pool. We therefore compare two protocols:
  • Shared Stage-1 pool. Stage-1 fine-tuning, $\mu$ estimation, and CD-T discovery inputs are all drawn from the same Stage-1 training pool.
  • Held-out discovery subset. The Stage-1 training pool is partitioned into disjoint subsets: one subset is used for Stage-1 fine-tuning, while the other is used for $\mu$ estimation and CD-T discovery inputs.

### v2

@@ D.2 Selection-ratio sweep for circuit construction (2%, 5%, 10%)
D.3 Shared vs. Held-out Competence Tuning Pool for Circuit Discovery (Iteration 0 Stability)
@@
Rather than constructing the baseline mean $\mu$ and the circuit-discovery inputs from fully disjoint data sources, we intentionally draw (i) the competence fine-tuning examples, (ii) the balanced-mean pool used to estimate $\mu$, and (iii) the circuit-discovery set from the same underlying training pool. This design choice is primarily practical and stability-driven: label-balanced mean estimation is constrained by the class distribution in low-resource splits, and using a separate, disjoint pool can reduce the availability of balanced samples and increase variance in $\mu$.
Importantly, this shared-pool choice is not merely about allowing overlap between the mean pool and discovery set; it is about ensuring that $\mu$ and the head-relevance estimates are computed from the same in-distribution pool as the checkpoint’s of competence tuning. This yields more stable circuits (especially at early iterations) and avoids degenerate selections that can arise when the baseline and discovery inputs are drawn from different pools. Concretely, ‘shared-pool’ refers to using Pool A for competence fine-tuning, $\mu$ estimation, and CD-T discovery; Pool B remains disjoint and is used only for mechanism-guided finetuning.
We observe in our low-resource setting that CD-T’s iteration-0 selection is more stable when circuit discovery is performed on the same pool used for competence fine-tuning, rather than on a held-out subset from that pool. We therefore compare two protocols:
  • Shared mechanism discovery pool. Competence fine-tuning, $\mu$ estimation, and CD-T discovery inputs are all drawn from the same mechanism discovery training pool.
  • Held-out discovery subset. The mechanism discovery pool is partitioned into disjoint subsets: one subset is used for competence fine-tuning, while the other is used for $\mu$ estimation and CD-T discovery inputs.